The height of a triangle is 2 m more than the base. The area of the triangle is 17.5m^2. Find the length of the base and the height of the triangle.
Answer by checkley77(12844)

H=B+2
A=BH/2
17.5=B*(B+2)/2
17.5=(B^2+2B)/2
B^2+2B=35
B^2+2B-35=0
(B+7)(B-5)=0
B-5=0
B=5 ANS. FOR THE BASE.
H=5+2=7 ANS FOR THE BASE
PROOF:
17.5=5*7/2
17.5=35/2
17.5=17.5